I need help on the following question:

(8r+3t)-|2-(t+3r)|

Answer by stanbon(75887)

(8r+3t)-|2-(t+3r)|
--------
2-(t+3r) = 8r+3t or 2-(t+3r) = -8r-3t
--------
2-t-3r = 8r+3t or 2-t-3r = -8r-3t
--------
4t+11r = 0 or 2t+5r = -2
===================================
Cheers,
Stan H.